# Sharp large deviation estimates for heavy-tailed extrema

### v1

@@ 1. Main result
This assumption covers a wide range of claim size distributions commonly used in actuarial science and finance, including Pareto, Student-$t$, Burr, stable distributions with index $\alpha<2$, and Loggamma distributions; see Embrechts et al. [4].
@@ 3. Proof of the main result
The following lemma describes the asymptotic behaviour of the normalizing sequence $(a_{n})$.
@@
In particular,
$a_{n}=n^{\frac{1}{\alpha}+o(1)}\qquad\text{as }n\to\infty.$ (3.3)
@@
We first note that $a_{n}\to\infty$. Indeed, if $(a_{n})$ were bounded, then $\bar{F}(x)=0$ for all sufficiently large $x$, implying that $X_{1}$ is almost surely bounded, which contradicts (3.4).
Fix $\varepsilon\in(0,1)$. By (3.4), there exists $x_{\varepsilon}>0$ such that for all $x\geq x_{\varepsilon}$,
@@
For $n$ large enough, $a_{n}-\varepsilon\geq x_{\varepsilon}$. By definition (1.1) of $a_{n}$,
@@
Using the lower bound in (3.5) together with (3.6), we obtain
@@
Moreover, by the definition of $a_{n}$ as an infimum, we have for every fixed $\varepsilon>0$,
@@
for all $n$ sufficiently large.
Applying the upper bound in (3.5) at $x=a_{n}-\varepsilon$, we obtain
@@
Letting $n\to\infty$ yields
@@
Since $\varepsilon>0$ is arbitrary, letting $\varepsilon\downarrow 0$ proves (3.2), and (3.3) follows immediately. ∎
@@
see [8, Proposition 1.15(b)]. This relation will be used below.
From (3.1), we write
@@
uniformly for $x\in[1,M]$.
Since $x\geq 1$, we have $t_{n}(x)\geq a_{n}\to\infty$. Moreover,
@@
For $u$ sufficiently close to $1$, the inequality $0\leq-\log u\leq 2(1-u)$ holds, and hence
$\log F(a_{n})\leq 2\bar{F}(a_{n})\leq\frac{2}{n},$
where we have used that $\bar{F}(a_{n})\leq 1/n$. It follows that
@@
We now study the fourth term
@@
Since the density $f$ is eventually non-increasing, the Von Mises condition (3.10) holds. Consequently, for $n$ large enough and all $x\in[1,M]$, we have
@@
Using Potter bounds iteratively, we may construct a sequence $\varepsilon_{n}\downarrow 0$ such that, for all $n$ sufficiently large and all $x\in[1,M]$,
$\frac{1-\varepsilon_{n}}{n}\,x^{-\frac{\log n}{\alpha}(\alpha+\varepsilon_{n})}\leq\bar{F}(t_{n}(x))\leq\frac{1+\varepsilon_{n}}{n}\,x^{-\frac{\log n}{\alpha}(\alpha-\varepsilon_{n})}.$ (3.12)
@@
$\frac{1}{\log n}\log f\bigl(t_{n}(x)\bigr)\leq\frac{1}{\log n}\log\!\left(\frac{\alpha+1}{t_{n}(x)}\right)+\frac{1}{\log n}\log\!\left(\frac{1+\varepsilon_{n}}{n}\right)-\left(1-\frac{\varepsilon_{n}}{\alpha}\right)\log x.$
@@
$\frac{1}{\log n}\log f\bigl(t_{n}(x)\bigr)\leq-\left(1+\frac{1}{\alpha}\right)-\left(1+\frac{1}{\alpha}\right)\log x+\frac{\varepsilon_{n}}{\alpha}\log M+o(1).$
Similarly, combining the lower bounds in (3.11) and (3.12), we obtain
$\frac{1}{\log n}\log f\bigl(t_{n}(x)\bigr)\geq-\left(1+\frac{1}{\alpha}\right)-\left(1+\frac{1}{\alpha}\right)\log x-\frac{\varepsilon_{n}}{\alpha}\log M+o(1).$
Together, these inequalities imply that
$\left|\frac{1}{\log n}\log f\bigl(t_{n}(x)\bigr)+\left(1+\frac{1}{\alpha}\right)+\left(1+\frac{1}{\alpha}\right)\log x\right|\leq\frac{\varepsilon_{n}}{\alpha}\log M+o(1),$
uniformly for $x\in[1,M]$. Since $\varepsilon_{n}\to 0$, we conclude that
@@
Collecting the limits of $T_{1}$, $T_{2}$, $T_{3}$, and $T_{4}$, we conclude that

### v2

@@ 1. Main result
This assumption covers a wide range of claim size distributions commonly used in actuarial science and finance, including the Pareto distribution, the Student-t distribution, the Burr distribution, stable distributions with index $\alpha<2$, and the log-gamma distribution; see Embrechts et al. [4].
@@ 3. Proof of the main result
The following lemma collects limits related to the normalizing sequence $(a_{n})_{n\in\mathbb{N}}$ that will be used in the proof of the main result.
@@
$\lim_{n\to\infty}\frac{\log\bar{F}(a_{n})}{\log n}=-1.$ (3.3)
@@
We first note that $a_{n}\to\infty$. Indeed, if $(a_{n})$ were bounded, we would have $\bar{F}(x)=0$ for all sufficiently large $x$, implying that the distribution has bounded support. This contradicts (3.4).
Fix $\varepsilon>0$. By (3.4), there exists $x_{\varepsilon}>0$ such that for all $x\geq x_{\varepsilon}$,
@@
Since $a_{n}\to\infty$, for $n$ large enough we have $a_{n}\geq x_{\varepsilon}$. By the definition of $a_{n}$,
@@
Using the lower bound in (3.5), we obtain
@@
Moreover, since $\bar{F}$ is non-increasing and $a_{n}$ is defined as an infimum, we have for every fixed $\varepsilon>0$,
@@
for all $n$ sufficiently large. For such $n$, we also have $a_{n}-\varepsilon\geq x_{\varepsilon}$, and applying the upper bound in (3.5) yields
@@
The second term on the right converges to zero as $n\to\infty$. Therefore,
@@
Letting $\varepsilon\downarrow 0$ proves (3.2).
Finally, it follows from (3.4) that
$\lim_{n\to\infty}\frac{\log\bar{F}(a_{n})}{\log a_{n}}=-\alpha.$
Consequently,
$\lim_{n\to\infty}\frac{\log\bar{F}(a_{n})}{\log n}$ $=\lim_{n\to\infty}\frac{\log\bar{F}(a_{n})}{\log a_{n}}\cdot\frac{\log a_{n}}{\log n}$
$=(-\alpha)\cdot\frac{1}{\alpha}=-1,$
where we have used (3.2). This proves (3.3). ∎
@@
see [8, Proposition 1.15(b)]. This asymptotic relation will be used below.
From (3.1), we may write
@@
with convergence uniform for $x\in[1,M]$.
Next, since $x\geq 1$, we have $t_{n}(x)\geq a_{n}\to\infty$. As $F$ is non-decreasing,
@@
For $u$ sufficiently close to $1$, the inequality $0\leq-\log u\leq 2(1-u)$ holds. Using $\bar{F}(a_{n})\leq 1/n$, we obtain
$\log F(a_{n})\leq 2\bar{F}(a_{n})\leq\frac{2}{n}.$
Consequently,
@@
We now turn to the fourth term,
@@
Since $f$ is eventually non-increasing, the Von Mises condition (3.10) implies that, for $n$ large enough and all $x\in[1,M]$,
@@
Applying iteratively the Potter bounds, we may construct a sequence $\varepsilon_{n}\downarrow 0$ such that, for all $n$ sufficiently large and all $x\in[1,M]$,
$(1-\varepsilon_{n})\,x^{-\frac{\log n}{\alpha}(\alpha+\varepsilon_{n})}\,\bar{F}(a_{n})\leq\bar{F}(t_{n}(x))\leq(1+\varepsilon_{n})\,x^{-\frac{\log n}{\alpha}(\alpha-\varepsilon_{n})}\,\bar{F}(a_{n}).$ (3.12)
@@
$\frac{1}{\log n}\log f\bigl(t_{n}(x)\bigr)$ $\leq\frac{\log(\alpha+1)}{\log n}-\frac{\log a_{n}}{\log n}-\frac{1}{\alpha}\log x$
$\quad+\frac{1}{\log n}\log\!\left(\frac{1+\varepsilon_{n}}{n}\right)-\left(1-\frac{\varepsilon_{n}}{\alpha}\right)\log x+\frac{1}{\log n}\log\bar{F}(a_{n}).$
@@
$\frac{1}{\log n}\log f\bigl(t_{n}(x)\bigr)\leq-\left(1+\frac{1}{\alpha}\right)-\left(1+\frac{1}{\alpha}\right)\log x+o(1),$
where the $o(1)$ term does not depend on $x$.
Similarly, combining the lower bounds in (3.11) and (3.12), we obtain that, uniformly for $x\in[1,M]$,
$\frac{1}{\log n}\log f\bigl(t_{n}(x)\bigr)\geq-\left(1+\frac{1}{\alpha}\right)-\left(1+\frac{1}{\alpha}\right)\log x+o(1).$
Together, these bounds imply
@@
Collecting the limits of $T_{1}$, $T_{2}$, $T_{3}$, and $T_{4}$, we observe that the constant terms cancel and the remaining contributions combine to give